A stone is thrown upward with an initial velocity of 50 𝑓𝑑/𝑠. What will its maximum height be? When will it strike the ground? Where will it be in 1.25𝑠?


1
Expert's answer
2021-09-25T13:09:33-0400

Given:

"v_0=50\:\rm ft/s"

"g=32.2\:\rm ft/s^2"


The maximum height reached by the stone

"h_{\max}=\frac{v_0^2}{2g}=\frac{50^2}{2*32.2}=38.8\:\rm ft"

The time of the motion of a stone

"t=2*\frac{v_0}{g}=2*\frac{50}{32.2}=3.1\:\rm s"

At the instant "t=1.25\:\rm s" the position of a stone

"h=v_0t-gt^2/2\\
=50*1.25-32.2*1.25^2/2=37.3\:\rm ft"